OpenStudy (anonymous):

What is the value of the sqrt(2+sqrt(2+sqrt(2+...to infinity?

OpenStudy (solomonzelman):

I would say it's approximately equal to square root of 2

OpenStudy (anonymous):

\[suppose~ x=\sqrt{2+\sqrt{2+\sqrt{2+\sqrt{2....}}}}\] \[Hence~x=\sqrt{2+x},x^2=2+x,x^2-x-2=0,x^2-2x+x-2=0,\] x(x-2)+1(x-2)=0 (x-2)(x+1)=0 x=2 , x=-1 (rejected, because sum of positive quantity cannot be negative. ,